OpenStudy (anonymous):

HELP..What is the y-value of sin(x) when x= -360°?

OpenStudy (zzr0ck3r):

0, 360, -360, 720, all get you to the same spot on the unit circle right?

OpenStudy (zzr0ck3r):

i.e. \(\sin(0) = \sin(360)=\sin(-360)=\sin(720)=........\) So the question reduces to what is \(\sin(0)= \ ?\)

OpenStudy (anonymous):

I dont get it